OpenStudy (highschoolmom2010):

a triangle has the sides 10,15,7. is it acute, obtuse or right? ANSWER CHECK

OpenStudy (highschoolmom2010):

\[10^2+15^2=7^2\] \[100+225=49\] \[325\neq49\] \[325>49\] it is acute correct?

OpenStudy (ybarrap):

@highschoolmom2010 , you are correct!

OpenStudy (ybarrap):

The hypotenuse is smaller than it would be if it were a right triangle, which is what you did. So smaller than 90 degrees means acute.